### v1

@@ 1 INTRODUCTION
Flexible yet robust architecture is a common feature of functional complex and adaptive systems [1]. Living systems across length- and timescales make use of a modular approach enabled by specific shared constraints that ensure proper functioning of the organism, while also deconstraining (i.e., conferring flexibilities upon) other aspects of the systems design without introducing risk of catastrophic consequences  [2]. This keeps the system functioning, while remaining adaptive and evolvable. Identification of these constraints provides a complementary framework to prevailing bottom-up mechanistic approaches and serves to directly probe core aspects of the how the system operates [3]. But in practice, how can we set about identifying these “constraints that deconstrain”? A promising approach is to design physics-quality, high-precision experiments to directly record the relevant dynamical behaviors, then sift the data for patterns of temporal organization that may provide helpful clues. However, to pursue this strategy, both technological and conceptual challenges must be overcome. To illustrate, we turn to the phenomenology of homeostasis, the self-regulating process wherein a complex system maintains its internal stability by holding key physiological variables constant within a desired tolerance [4]. This is similar to mechanical self-regulating control systems (such as thermostats); indeed, the mathematical framework describing feedback control in each process shares a common origin [5]. However, any characterization of organismal homeostases from a systems design point of view must distinguish between two opposing paradigms: passive “elastic adaptation” characterized by reflexive responses to present conditions with no memory of the past (as in Ashby’s Homeostat [6]), versus active “plastic adaptation” involving reflective responses achieved by controlled memory integrators operating on information stored from past experiences (as in Shannon’s “Maze Solving Machine” [7]). In contrast, most extant theoretical models of homeostasis have approached the problem using a (quasi-)deterministic framework built on the assumption of a defined “setpoint” value of the state variable; this approach captures the behaviors of the mythical “average” cell, yet cannot generate a quantitative picture of the stochastic homeostasis of a realistic individual cell, and thus falls short of proving mechanistic insights. To understand the origins of this conundrum, and motivate the pathway to move beyond it, we briefly review the history of bacterial growth physiology and the feedback between experimental and theoretical advances.
@@
In the 1940s, novel techniques for continuous culture of bacterial colonies such as the Chemostat [8] permitted researchers to maintain populations at a constant rate of numbers growth for indefinite periods of time [9]. These studies of asynchronous bacterial populations in steady state—typically referred to as “balanced growth” [10]—enabled the first reproducible and quantitative growth studies, leading to the seminal discoveries that total population growth is linearly proportional to the initial concentration of the rate-limiting nutrient [11] and that average cell size is exponentially proportional to average nutrient-imposed growth rate [12]. Shortly thereafter, researchers quantified the growth curves of single cells using time-lapse microscopy of small populations of cells growing over 3–4 generations on an agar surface sandwiched inside a sealed chamber [13]. Technical limitations prevented determination of the precise mode of growth, but observations were consistent with the hypothesis that cells grow exponentially at roughly the same rate as the whole population [14]. The next major experimental refinement occurred with the introduction of the “Baby Machine” [15], a device for creating a population of cells synchronized according to the most recent cell division. This method permitted the first studies into the timing of events relative to the cell cycle, demonstrating that the onset of DNA replication varies with cell growth rate [16] and occurs when the cell mass per chromosome reaches a (constant) critical value, thus explaining the previously observed relationship between average cell size and growth rate [17].
@@
For decades, however, high-precision single-cell measurements were accessible only through snapshots of cells drawn from steady-state populations after being “fixed” (i.e., preserved in a life-like state for imaging). This approach yielded important results, including the observation that mean-rescaled (asynchronous) cell size distributions obtained from various growth conditions undergo a scaling collapse [18]. However, single-cell growth studies capable of obtaining high-precision, dynamic information did not become widespread until the early 2000s, when researchers developed protocols to more readily and reproducibly integrate growth of cells on flat agarose pads with time-lapse microscopy to obtain movies of individual living cells, allowing for the study of intracellular spatiotemporal organization and morphology [19], single-cell growth in dynamic environments [20], and cell size homeostasis [21]. Due to its ease of use, this technique remains indispensable for high-throughput screens linking genotype to phenotype, permitting assays across thousands of combinations of bacterial mutants and growth conditions to link molecular-level details to cell size and growth [22]. However, exponential numbers growth of cells limits the duration and interpretability of such experiments [23]—especially problematic when growth itself is the process of interest. The introduction of microfluidic technology sought to address this problem by enabling growth of microcolonies in a precisely controlled chemostat-like environment compatible with long-term imaging [24]. A variety of approaches have been explored [25, 26, 27], but the design that has proved the most popular, known as the “Mother Machine”, cultivates thousands of cell lineages inside microscope channels open to flowing growth medium at one end and closed at the other; although daughter cells eventually exit the channel, trapped mother cells may be observed for over 100 generations [28]. This technology has been applied to studies of single-cell aging (or lack thereof) [28, 29], cell-fate decision making [30], and cell size homeostasis [31, 32, 33, 34] at steady state, as well as physiological growth [35] and adaptation [36] in time-varying environments. A complementary approach is the SChemostat [37] design, which uses controllable surface adhesion to generate a stable population of cells of a defined density chosen to eliminate neighbor–neighbor contacts and ensure that all cells experience an identical flow environment. By removing potential confounding factors such as mechanical confinement or variable microenvironments [38], this permits observation of a statistically identical set of non-interacting cells under precisely controlled and tunable environmental conditions for indefinite periods of time, a prerequisite for detailed studies of stochastic growth and division over many generations.
@@ 2.1.2 Experimental evidence
But does cell size, which serves as a proxy for cell mass, in fact grow exponentially between division events? Providing an unequivocal answer through direct experimentation has proved challenging; whereas population growth under balanced conditions is readily measured to be exponential and explained as the result of cell doubling at a constant rate [11], establishing the temporal law governing cell size growth remained an open problem for decades. The dominant paradigms considered were linear growth—the simplest possible variation with time—and exponential growth—a “natural” choice (see the discussion regarding the Hinshelwood cycle above), with suggestive but inconclusive evidence from early single-cell growth studies [13]. The difficulty in answering this question arises from the necessity of experimentally resolving the difference between a linear and an exponential curve over a $\sim$2-fold change in cell size between the initial and final points of an individual generation of the bacterial growth curve (requiring much better than $4\%$ resolution [37]). This is difficult owing to the substantial noise present in bacterial growth curves, even though they clearly show a curvature (i.e., are non-linear) to the naked eye [13, 37]. High-precision SChemostat experiments on large numbers of statistically identical, non-interacting cells allowed this distinction to be resolved via statistical averaging (Fig. 1a–d), establishing beyond reasonable doubt that the pre-divisional C. crescentus stalked cell grows exponentially with time under balanced growth conditions [37]. Based on such measurements of single-cell size as well as cell mass [51], it is now considered a widely accepted fact that sizes of pre-divisional bacterial cells, on average, increase exponentially under balanced growth conditions.
@@ 2.1.3 Constraining Motif: The Stochastic Hinshelwood Cycle
The scaling law governing the intragenerational temporal evolution of cell sizes points to the emergence of an elegant constraining motif from the otherwise confounding complexity and system- and condition-specific details of cellular metabolic networks: the Stochastic Hinshelwood Cycle (SHC) [52]. First, following the discussion of the deterministic Hinshelwood Cycle above, we see that robustness of intergenerational continuity of exponential growth is satisfied by a noise-free Hinshelwood Cycle scheme of reactions. Inclusion of stochasticity through the most prevalent phenomenological model of stochastic exponential growth, geometric Brownian motion with linear multiplicative noise (which forms the basis of the widely used Black-Scholes equation in economics), predicts runaway noise in cell sizes [52]. This is manifestly in contradiction with observations. Instead, a generalization of the Hinshelwood cycle that includes stochasticity was found to represent stochastic exponential growth with square-root multiplicative noise (Fig. 1e) [52]; it belongs to a continuous family of well-behaved models of stochastic exponential growth with fractional power multiplicative noise [53]. The prediction of square-root multiplicative noise was directly verified from high-precision data [37]. Another prediction of the SHC model is that mean-rescaled cell sizes follow the same distribution at all times throughout the cell cycle [52]. This prediction has been validated by a spectacular multidimensional scaling collapse, across generation times and temperatures, of cell sizes from high-precision data (Fig. 1f,g) [37]. Yet the question remains: How does the cartoonishly simple Stochastic Hinshelwood Cycle relate to actual complex biochemical networks governing cell growth in system and condition specific ways? We provide the answer in Section (2.2.4).
@@ 2.2.1 Implications of the Stochastic Hinshelwood Cycle motif
Overall population dynamics in a given growth condition are primarily determined by two timescales: that of cell size growth and that of population number growth. The latter is determined by the interdivision time interval (henceforth referred to simply as the “division time”) [54]. Since the SHC predicts that all constituent species grow at the same exponential growth rate in balanced growth, this naturally establishes the timescale of cell size growth as the inverse of this exponential growth rate. This rate is determined by the geometric mean of the reaction rates of the dominant cycle in the SHC (Fig. 1h–j). Hence, if on varying the ambient temperature the individual rates follow the Arrhenius law with activation energy barriers characteristic of typical enzyme reactions, then so must the composite cell size growth rate! Imposing the constraint of cell size homeostasis couples these two timescales, implying that a single cellular unit of time governs stochastic growth and division of individual cells in a given condition. We address the specifics of this coupling in Section (3.1.7), but intuitively: if the population numbers grow faster (slower) than the total cell size, the average cell size will decrease (increase).
@@ 2.2.2 Experimental validation of a single cellular unit of time
The expectation that a single cellular unit of time governs stochastic growth and division of individual cells is spectacularly borne out by experimental data (Fig. 2a–g). The coupling of the timescales of size and number growth is evidenced from the observation that the product of the mean growth rate and the mean division time is remarkably invariant across different temperatures [37]. Furthermore, the mean division time changes with temperature according to the Arrhenius law with an activation energy barrier of approximately $12.9$ kcal/mol [37] (within the range of temperatures of physiological relevance). This is of the same order of magnitude as the activation energy of a typical enzyme-catalyzed reaction in the cell, providing compelling validation for the SHC model. This Arrhenius-like energy scale of single-cell growth rates matches remarkably well with measurements of population growth rates of multiple species as a function of temperature [55, 56, 57].
@@ 2.2.3 Reconciling the mythical average cell with the stochastic single cell
This scaling was demonstrated in C. crescentus cells through high-precision SChemostat experiments at different temperatures, establishing that the division time distributions, which can vary substantially across physiologically accessible temperatures, nevertheless collapse to the same distribution when rescaled by their respective mean values (Fig. 3a,b) [37]! We label this “mean-rescaling”. Since the population age distribution is determined by the dynamics of cell division, population age distributions across a range of temperatures also undergo mean-rescaling (Fig. 3c) [54]. Mean-rescaling of both size and division time distributions across different conditions have also been observed in E. coli [61, 62], hinting at the universality of a characteristic timescale determining growth and division dynamics across different bacterial species. Scaling has also been shown for the distributions of protein copy numbers measured in populations under a wide range of conditions in E. coli by first subtracting the mean and then dividing by the standard deviation [63].
@@ 3.1.2 Experimental evidence for stochastic intergenerational homeostasis
It has only recently been rigorously established via high-precision experiments that cell sizes are maintained over many generations [64]. Since a cell’s size changes over the course of a generation, a precise analysis requires the choice of a representative size within each generation, that can be compared to a comparable size in other generations. We choose this representative size to be the initial cell size of a given generation, and consider its intergenerational dynamics (Fig. 4a–c). High-precision long-term experiments using the SChemostat have established that not only does the population size distribution remain invariant in balanced growth conditions (Fig. 4d,e), but also that subpopulations characterized by different sizes in the starting (first-observed) generations all converge to the same homeostatic distribution of cell sizes after sufficiently many generations have elapsed (Fig. 4f) [64]. This nontrivial probabilistic ergodic behavior conclusively demonstrates stochastic intergenerational homeostasis of cell size [64].

### v2

@@ 1 INTRODUCTION
Flexible yet robust architecture is a common feature of functional complex and adaptive systems [1]. Living systems across length- and timescales make use of a modular approach enabled by specific shared constraints that ensure proper functioning of the organism, while also deconstraining (i.e., conferring flexibilities upon) other aspects of the systems design without introducing risk of catastrophic consequences  [2]. This keeps the system functioning, while remaining adaptive and evolvable. Identification of these constraints provides a complementary framework to prevailing bottom-up mechanistic approaches and serves to directly probe core aspects of how the system operates [3]. But in practice, how can we set about identifying these “constraints that deconstrain”? A promising approach is to design high-throughput, high-precision experiments to directly record the relevant dynamical behaviors, then sift the data for patterns of temporal organization that may provide helpful clues. However, to pursue this strategy, both technological and conceptual challenges must be overcome. To illustrate, we turn to the phenomenology of homeostasis, the self-regulating process wherein a complex system maintains its internal stability by holding key physiological variables constant within a desired tolerance [4]. This is similar to mechanical self-regulating control systems (such as thermostats); indeed, the mathematical framework describing feedback control in each process shares a common origin [5]. However, any characterization of organismal homeostases from a systems design point of view must distinguish between two opposing paradigms: passive “elastic adaptation” characterized by reflexive responses to present conditions with no memory of the past (as in Ashby’s Homeostat [6]), versus active “plastic adaptation” involving reflective responses achieved by controlled memory integrators operating on information stored from past experiences (as in Shannon’s “Maze Solving Machine” [7]). In contrast, most extant theoretical models of homeostasis have approached the problem using a (quasi-)deterministic framework built on the assumption of a defined “setpoint” value of the state variable; this approach captures the behaviors of the mythical “average” cell, yet cannot generate a quantitative picture of the stochastic homeostasis of a realistic individual cell, and thus falls short of proving mechanistic insights. To understand the origins of this conundrum, and motivate the pathway to move beyond it, we briefly review the history of bacterial growth physiology and the feedback between experimental and theoretical advances.
@@
In the 1940s, novel techniques for continuous culture of bacterial colonies such as the Chemostat [8] permitted researchers to maintain populations at a constant rate of growth for indefinite periods of time [9]. These studies of asynchronous bacterial populations in steady state—typically referred to as “balanced growth” [10]—enabled the first reproducible and quantitative growth studies, leading to the seminal discoveries that total population growth is linearly proportional to the initial concentration of the rate-limiting nutrient [11] and that average cell size is exponentially proportional to average nutrient-imposed growth rate [12]. Shortly thereafter, researchers quantified the growth curves of single cells using time-lapse microscopy of small populations of cells growing over 3–4 generations on an agar surface sandwiched inside a sealed chamber [13]. Technical limitations prevented determination of the precise mode of growth, but observations were consistent with the hypothesis that cells grow exponentially at roughly the same rate as the whole population [14]. The next major experimental refinement occurred with the introduction of the “Baby Machine” [15], a device for creating a population of cells synchronized according to the most recent cell division. This method permitted the first studies into the timing of events relative to the cell cycle, demonstrating that the onset of DNA replication varies with cell growth rate [16] and occurs when the cell mass per chromosome reaches a (constant) critical value, thus explaining the previously observed relationship between average cell size and growth rate [17].
@@
For decades, however, high-precision single-cell measurements were accessible only through snapshots of cells drawn from steady-state populations after being “fixed” (i.e., preserved in a life-like state for imaging). This approach yielded important results, including the observation that mean-rescaled (asynchronous) cell size distributions obtained from various growth conditions undergo a scaling collapse [18]. However, single-cell growth studies capable of obtaining high-precision, dynamic information did not become widespread until the early 2000s, when researchers developed protocols to more readily and reproducibly integrate growth of cells on flat agarose pads with time-lapse microscopy to obtain movies of individual living cells, allowing for the study of intracellular spatiotemporal organization and morphology [19], single-cell growth in dynamic environments [20], and cell size homeostasis [21]. Due to its ease of use, this technique remains indispensable for high-throughput screens linking genotype to phenotype, permitting assays across thousands of combinations of bacterial mutants and growth conditions to link molecular-level details to cell size and growth [22]. However, exponential growth of cell numbers limits the duration and interpretability of such experiments [23]—especially problematic when growth itself is the process of interest. The introduction of microfluidic technology sought to address this problem by enabling growth of microcolonies in a precisely controlled chemostat-like environment compatible with long-term imaging [24]. A variety of approaches have been explored [25, 26, 27], but the design that has proved the most popular, known as the “Mother Machine”, cultivates thousands of cell lineages inside microscope channels open to flowing growth medium at one end and closed at the other; although daughter cells eventually exit the channel, trapped mother cells may be observed for over 100 generations [28]. This technology has been applied to studies of single-cell aging (or lack thereof) [28, 29], cell-fate decision making [30], and cell size homeostasis [31, 32, 33, 34] at steady state, as well as physiological growth [35] and adaptation [36] in time-varying environments. A complementary approach is the SChemostat [37] design, which uses controllable surface adhesion to generate a stable population of cells of a defined density chosen to eliminate neighbor–neighbor contacts and ensure that all cells experience an identical flow environment. By removing potential confounding factors such as mechanical confinement or variable microenvironments [38], this permits observation of a statistically identical set of non-interacting cells under precisely controlled and tunable environmental conditions for indefinite periods of time, a prerequisite for detailed studies of stochastic growth and division over many generations.
@@ 2.1.2 Experimental evidence
But does cell size, which serves as a proxy for cell mass, in fact grow exponentially between division events? Providing an unequivocal answer through direct experimentation has proved challenging; whereas population growth under balanced conditions is readily measured to be exponential and explained as the result of cell doubling at a constant rate [11], establishing the temporal law governing cell size growth remained an open problem for decades. The dominant paradigms considered were linear growth—the simplest possible variation with time—and exponential growth—a “natural” choice (see the discussion regarding the Hinshelwood cycle above), with suggestive but inconclusive evidence from early single-cell growth studies [13]. The difficulty in answering this question arises from the necessity of experimentally resolving the difference between a linear and an exponential curve over a $\sim$2-fold change in cell size between the initial and final points of an individual generation of the bacterial growth curve (requiring much better than $4\%$ resolution [37]). This is difficult owing to the substantial noise present in bacterial growth curves, even though they clearly show a curvature (i.e., are non-linear) to the naked eye [13, 37]. High-precision SChemostat experiments on large numbers of statistically identical, non-interacting cells, consisting of multigenerational trains of growth curves corresponding to each individual cell (Fig. 1a), which can be further decomposed into growth curves corresponding to each generation of each cell (Fig. 1b), allowed this distinction to be resolved via statistical averaging; this established beyond reasonable doubt that the pre-divisional C. crescentus stalked cell grows exponentially with time under balanced growth conditions [37]. Based on such measurements of single-cell size as well as cell mass [51], it is now considered a widely accepted fact that sizes of pre-divisional bacterial cells, on average, increase exponentially under balanced growth conditions.
@@ 2.1.3 Constraining motif: The Stochastic Hinshelwood Cycle
The scaling law governing the intragenerational temporal evolution of cell sizes points to the emergence of an elegant constraining motif from the otherwise confounding complexity and system- and condition-specific details of cellular metabolic networks: the Stochastic Hinshelwood Cycle (SHC) [52]. First, following the discussion of the deterministic Hinshelwood Cycle above, we see that robustness of intergenerational continuity of exponential growth is satisfied by a noise-free Hinshelwood Cycle scheme of reactions. Inclusion of stochasticity through the most prevalent phenomenological model of stochastic exponential growth, geometric Brownian motion with linear multiplicative noise (which forms the basis of the widely used Black-Scholes equation in economics), predicts runaway noise in cell sizes [52]. This is manifestly in contradiction with observations. Instead, a generalization of the Hinshelwood cycle that includes stochasticity was found to represent stochastic exponential growth with square-root multiplicative noise [52]; it belongs to a continuous family of well-behaved models of stochastic exponential growth with fractional power multiplicative noise [53]. The prediction of square-root multiplicative noise was directly verified from high-precision data [37]. Another prediction of the SHC model is that mean-rescaled cell sizes follow the same distribution at all times throughout the cell cycle [52]. As an illustrative example, consider a simplified SHC consisting of three species (Fig. 1e). Simulations of of the time evolution of the copy numbers of each species demonstrate that all species grow at the same exponential rate at steady state (Fig. 1f); furthermore, the distributions of copy numbers of each species, upon rescaling by their respective mean values, are identical (Fig. 1g). This prediction has been validated by a spectacular multidimensional scaling collapse, across generation times and temperatures, of cell sizes from high-precision data [37]. Thus, the experimentally measured distributions of cell sizes as a function of cell age (defined as time since last division) are distinct for a given cell age (Fig. 1c); however, after rescaling all distributions by their respective mean values, they collapse onto a single distribution (Fig. 1d). Yet the question remains: How does the cartoonishly simple Stochastic Hinshelwood Cycle relate to actual complex biochemical networks governing cell growth in system and condition specific ways? We provide the answer in Section (2.2.4).
@@ 2.2.1 Implications of the Stochastic Hinshelwood Cycle motif
Overall population dynamics in a given growth condition are primarily determined by two timescales: that of cell size growth and that of population number growth. The latter is determined by the interdivision time interval (henceforth referred to simply as the “division time”) [54]. Since the SHC predicts that all constituent species grow at the same exponential growth rate in balanced growth, this naturally establishes the timescale of cell size growth as the inverse of this exponential growth rate. This rate is determined by the geometric mean of the reaction rates of the dominant cycle in the SHC. All rates in the SHC are expected to scale by the same factor as temperature varies (provided it remains in the range where balanced growth conditions hold) (Fig. 1h). Although the connections within the SHC may change between balanced growth conditions (Fig. 1i), the overall growth rate can be determined by decomposing the network into constitutive loops and identifying the loop with the highest geometric mean of individual rates (Fig. 1j). Hence, if on varying the ambient temperature the individual rates follow the Arrhenius law with activation energy barriers characteristic of typical enzyme reactions, then so must the composite cell size growth rate! Imposing the constraint of cell size homeostasis couples these two timescales, implying that a single cellular unit of time governs stochastic growth and division of individual cells in a given condition. We address the specifics of this coupling in Section (3.1.7), but intuitively: if the population numbers grow faster (slower) than the total cell size, the average cell size will decrease (increase).
@@ 2.2.2 Experimental validation of a single cellular unit of time
The expectation that a single cellular unit of time governs stochastic growth and division of individual cells is spectacularly borne out by experimental data. Cell size trajectories at temperatures spanning the physiologically relevant range for C. crescentus, from 34 C (Fig. 2a) to 17 C (Fig. 2b), show markedly different dynamics. To compare these data, note that each growth curve may be characterized by the initial size ($a_{i}$), final size ($a_{f}$), growth rate ($k$), and division time ($\tau$) (Fig. 2c). C. crescentus has a complex dimorphic life cycle involving asymmetric division into a stalked and a swarmer cell, the latter of which must differentiate into a stalked cell before it is capable of replication (Fig. 2d). However, in the SChemostat, swarmer cells are removed immediately after division; thus, out of all possible future lineages of a given cell (Fig. 2e), only the single lineage consisting exclusively of stalked cells is observed, and the quiescence timescale need not be considered. For a given condition, the timescale governing growth and division may thus be represented by the distribution of division times (Fig. 2f) and the distribution of cell ages (Fig. 2g). The coupling of the timescales of size and number growth is evidenced from the observation that the product of the mean growth rate and the mean division time is remarkably invariant across different temperatures [37]. Furthermore, the mean division time changes with temperature according to the Arrhenius law with an activation energy barrier of approximately $12.9$ kcal/mol [37] (within the range of temperatures of physiological relevance). This is of the same order of magnitude as the activation energy of a typical enzyme-catalyzed reaction in the cell, providing compelling validation for the SHC model. This Arrhenius-like energy scale of single-cell growth rates matches remarkably well with measurements of population growth rates of multiple species as a function of temperature [55, 56, 57].
@@ 2.2.3 Reconciling the mythical average cell with the stochastic single cell
This scaling was demonstrated in C. crescentus cells through high-precision SChemostat experiments at different temperatures, establishing that the division time distributions, which can vary substantially across physiologically accessible temperatures (Fig. 3a), nevertheless collapse to the same distribution when rescaled by their respective mean values (Fig. 3b) [37]! We label this “mean-rescaling”. Since the population age distribution is determined by the dynamics of cell division, population age distributions across a range of temperatures also undergo mean-rescaling (Fig. 3c) [54]. Mean-rescaling of both size and division time distributions across different conditions have also been observed in E. coli [61, 62], hinting at the universality of a characteristic timescale determining growth and division dynamics across different bacterial species. Scaling has also been shown for the distributions of protein copy numbers measured in populations under a wide range of conditions in E. coli by first subtracting the mean and then dividing by the standard deviation [63].
@@ 3.1.2 Experimental evidence for stochastic intergenerational homeostasis
It has only recently been rigorously established via high-precision experiments that cell sizes are maintained over many generations [64]. Since a cell’s size changes over the course of a generation, a precise analysis requires the choice of a representative size within each generation, that can be compared to a comparable size in other generations. We choose this representative size to be the initial cell size of a given generation (Fig. 4a), and consider its intergenerational dynamics. Quasi-deterministic models of homeostasis within the sizer–adder–timer paradigm envision the dynamics of initial cell sizes as an orderly regression to the population mean (Fig. 4b), but this does not reflect the trajectories of actual cells measured from high-precision, long-term experiments using the SChemostat (Fig. 4c). To capture these stochastic dynamics, consider the conditional distributions of the next generation’s initial sizes, conditional on the current generation’s initial size (Fig. 4d). When the conditional distributions for different values of the current generation’s initial size are rescaled by their respective mean values, they collapse onto a single distribution (Fig. 4e). In addition, the conditional distributions of the initial sizes after $n$ generations, conditional on the current generation’s initial size, all converge to the steady-state distribution of initial sizes for $n\gg 1$ (Fig. 4f). Thus, not only does the population size distribution remain invariant in balanced growth conditions, but subpopulations characterized by different sizes in the starting (first-observed) generation all converge to the same homeostatic distribution of cell sizes after sufficiently many generations have elapsed. This nontrivial probabilistic ergodic behavior conclusively demonstrates stochastic intergenerational homeostasis of cell size [64].